For the following reaction (which is exothermic in
Answer #1

2SO2 + O2 <-------> 2SO3 + heat (exothermic)

a. The equilibrium will shift to the right in order to consume the O2 added

b. You have to use heat as a product or reagent, in these case the reaction is exothermic so it releases heat, it will be a "product". If we add heat the reaction will shift to the left in order to consume the heat.

c. Opposite of b

d. We are taking away a product so the reaction will shift to the left in order to produce the product.

e. I´m not sure about this one. If the H2 doesn´t react with anyone then the equilibrium position won´t change.

--------------------------------------------------------------------------------------------

a. yes, the precipitate will be PbI2

b. yes, the precipitate will be AgCl

c. no precipitate

d. yes, the precipitate will be CaSO4

e. no precipitate

----------------------------------------------------------------------------------

%m/m= mass solute/mass solution x 100

The density of water is 1g/mL so, 325mL are the same as 325g

%m/m= 85.635g/325g+85.635g x100=20.9%

---------------------------------------------------------------------------------

mol NaCl= 36.063g/58g/mol= 0.622 mol

M= mol/V(L)= 0.622mol/1.25L= 0.497M

-----------------------------------------------------------------------------------

V= mol/M= 0.3mol/0.115M= 2.61 L

-----------------------------------------------------------------------------------

Ci x Vi= Cf x Vf

Cf= 12.4M x 8.5x10-3L/0.35L = 0.301M

-------------------------------------------------------------------

3H2SO4 + 2Fe(OH)3 -----> Fe2(SO4)3 + 6H2O

[OH-]x[H+]= 1x10-14M ----> [OH-]= 2.58x10-12M

pH= -log[H+]= -log 3.88x10-3M= 2.41

pOH= -log[OH-]= -log 2.58x10-12M= 11.59